OpenStudy (anonymous):

what is 0.194 with 94 repeating into a fraction? will medal

OpenStudy (campbell_st):

let x = 0.19494949494...... multiply the decimal by 1000 (equation 1) and by 10 (equation 2) find the difference of the 2 equations equation 1 - equation 2 get an answer then solve for x

OpenStudy (anonymous):

what do you mean equation 1 - equation 2?

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

@campbell_st can u explain again?

OpenStudy (campbell_st):

ok...x = 0.194949494..... so mutiply by a value that only has the recurring decimal to the left of the decimal point. you will need 2 equations... and you and normally multiplying by multiples of 10. then then you have 2 equations subtract the smaller from the larger 1000x = 194.9494949494 equation 1 - 10x = 1.9494949494 equation 2 ----------------------- get the answer... and you should see the recurring decimal disappears... then its a case of just solving for x

OpenStudy (anonymous):

so 193 over 999?

OpenStudy (campbell_st):

you will get a fraction as the answer... that may need to be simplified. as an example 0.2222222222 10x = 2.22222222 - x = 0.22222222 ----------------- 9x = 2 then x = 2/9 so 0.222222 as a fraction is 2/9 hope it makes sense

OpenStudy (campbell_st):

no, its not 999 its 990...

OpenStudy (campbell_st):

but otherwise your solution is great

OpenStudy (anonymous):

oh so 193 over 990?

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

thanks

OpenStudy (campbell_st):

these questions all work the same... it just a case of identifying the multiple of 10 to use